Father's age is 5 times the sum of the age of his children.In 8 years, the father's age will be twice the sum of the ages of these 2 children.Find his present age,
Answer by richwmiller(17219)

f=5*(a+b),
f+8=2*(a+8+b+8)
b = 8-a, f = 40